### v1

# thanks: These authors contributed equally to this workthanks: These authors contributed equally to this work

Abstract
Driven-dissipative many-body system supports nontrivial quantum phases absent in equilibrium. As a prominent example, the interplay between coherent driving and collective dissipation can lead to a dynamical quantum phase that spontaneously breaks time-translation symmetry. This so-called boundary time crystal (BTC) is fragile in the presence of local dissipation, which inevitably relaxes the system to a stationary state. In this work, we demonstrate a robust BTC that is intrinsically induced by local dissipation. We provide extensive numerical evidences to support existence of the BTC and study its behaviors in different regimes. In particular, with decreasing interaction range, we identify a transition from classical limit cycles to quantum BTCs featuring sizable spatial correlations. Our studies significantly broaden the scope of nonequilibrium phases and shed new light on experimental search for dynamical quantum matter.
preprint: APS/123-QED
Introduction.—Spontaneous symmetry breaking in the temporal dimension manifests itself in a nontrivial quantum phase: the continuous time crystal (CTC) [1]. In the CTC phase, a small perturbation brings the system away from the stationary state and triggers a persistent, ordered oscillation of local observables [2, 3, 4, 5]. While existence of this intriguing phase in equilibrium has been questioned by a series of studies [6, 7, 8], it has been predicted that CTC can be realized in a driven-dissipative setup [9, 10, 11, 12, 13, 14, 15, 16, 17]. This possibility is highlighted by a number of recent experiments observing self-sustained oscillations, such as in ultracold atoms [18, 19, 20], thermal Rydberg gases [21, 22, 23, 24, 25, 26], and solid-state materials [27, 28].
The dissipative version of the CTC is also called a boundary time crystal (BTC), considering the system as a boundary in contact with a bulk environment [29]. The prototype model of the BTC considers a system of spin-1/2 particles subjected to a superradiant decay [30], which can be induced by a propagating bath mode in the Markovian regime [see Fig. 1(a)]. However, a realistic setup inevitably carries local decay channels, which immediately destroy the established oscillation pattern in the long-time limit [31, 32], as shown in Fig. 1(b).
Inspired by recent experiments in Rydberg atomic ensembles [22, 23, 24] we study in this Letter a one-dimensional (1D) lattice model, where long-range interacting spin-1 particles experience strictly local decay [see Fig. 1(c)]. While local dissipations usually erase quantum correlations [33], we show that, remarkably, they can induce a BTC phase here [see Fig. 1(d)]. At the mean-field level, this BTC corresponds to a classical limit cycle, which originates from competition between distinct spin coherences. We perform extensive numerical simulations to investigate the quantum nature and origin of the BTC. In the case of all-to-all interactions, we show that the Liouvillian spectrum exhibit gapless and ordered features, while persistent oscillation of the two-time correlation function unambiguously reveals a spontaneous time-translation symmetry breaking. For a power-law interaction $\sim r^{-\alpha}$, our analysis reveals a critical $\alpha_{c}$, above which the interaction is too short-ranged to support the BTC. Interestingly, $\alpha_{c}$ is larger than the dimension of the system, giving rise to a quantum BTC without a classical correspondence in the regime $1<\alpha<\alpha_{c}$. In a quantum BTC, considerable spatial correlations can be rapidly established from an uncorrelated initial state.
Figure 1: (a) and (b) show the schematic of the superradiance-based BTC and its mean-field evolution of the order parameter $m_{z}$, respectively. Here, an arbitrarily small local decay ($\gamma\neq 0$) can destroy the established oscillation. (c) and (d) show the proposed spin-1 BTC model and its mean-field evolution of the order parameter $s_{z}$, respectively. Here, the oscillation is robust against local dissipation and persists within a finite range of the decay rate $\gamma$.
Microscopic model.—We consider a spin-1 model in a 1D lattice [see Fig. 1(c)], described by the Hamiltonian
${H}=\frac{\Omega}{\sqrt{2}}\sum_{i}{s}_{i}^{x}-\Delta\sum_{i}{n}_{i}-E\sum_{i}{s}_{i}^{z}-\sum_{i<j}V_{ij}{n}_{i}{n}_{j},$ (1)
where the spin-1 operators are defined as $s_{i}^{x}=(|0\rangle_{i}\langle+|+|0\rangle_{i}\langle-|+\mathrm{H.c.})/\sqrt{2}$, $s_{i}^{z}=|+\rangle_{i}\langle+|-|-\rangle_{i}\langle-|$, and $n_{i}=(s_{i}^{z})^{2}=|+\rangle_{i}\langle+|+|-\rangle_{i}\langle-|$. In such a $\mathsf{V}$-type excitation scheme, the spin state |$\ket{0}$ is coherently coupled to states $\ket{+}$ and $\ket{-}$ (energetically separated by $E$) with a Rabi frequency $\Omega$ and a detuning $\Delta$ from the average energy level of the exited states. Meanwhile, spins in states $\ket{\pm}$ feature a power-law interaction $V_{ij}=C/|i-j|^{\alpha}$. In addition to coherent drivings and interactions, each spin in states $\ket{\pm}$ decays individually to the state |$\ket{0}$ in a Markovian manner with a rate $\gamma$. The system density matrix $\rho$ then evolves according to a Lindblad master equation $\partial_{t}\rho=\mathcal{L}(\rho)$, where the Liouvillian operator acts as
[…]
with $L_{j}^{\sigma}=\sqrt{\gamma}|0\rangle_{j}\langle\sigma|$ the local jump operator at site $j$. This model has a neutral-atom implementation, where |$\ket{0}$ denotes the ground state of an atom, and $\ket{\pm}$ represent two distinct Rydberg states [34, 35, 36], respectively.
Mean-field analysis.—We first treat the system at the mean-field level, in which we neglect the correlation between different sites and assume that local operators $o_{i}$ have a homogeneous expectation value, i.e., $\langle o_{i}o_{j}\rangle=\langle o_{i}\rangle\langle o_{j}\rangle$ for $j\neq k$ and $\langle{o}_{i}\rangle=\langle{o}_{j}\rangle$. This treatment will generate a set of nonlinear equations [37], whose nonlinear dynamics is solely determined by the collective interaction strength $\chi=\sum_{j}V_{ij}$.
We then perform stability analysis of the equations of motion [38] and study the long-time behavior of the system. First, when states $\ket{\pm}$ are degenerate ($E=0$), the system only has stable fixed points, corresponding to a stationary phase (SP) where observables rapidly decay to the steady-state values. Within the SP, by tuning the driving parameters $\Omega$ and $\Delta$, the system can undergo a first-order phase transition between a bistable phase and a trivial one. In fact, although states $\ket{\pm}$ decay independently, the mean-field equations at $E=0$ is equivalent to those of a spin-1/2 model, which is widely used to describe the optical bistability in Rydberg atomic ensembles [39, 40, 41, 42, 43].
As the energy separation $E$ increases, the competition between states $\ket{\pm}$ comes into play, which is quantified by the order parameter $s_{z}=\langle s_{i}^{z}\rangle$ measuring their population difference. Near the intermediate driving ($|\Delta|<\Omega$), the system supports a Hopf bifurcation at which the stability of a fixed point suddenly changes (see the Supplemental Material for details [37]). Figure 1(d) shows a typical case $E=4$, $\Delta=-1$, $\Omega=3$, $\chi=16$, $\gamma=1$ (used throughout this work), in which $s_{z}$ exhibits a persistent periodic oscillation, whose trajectory is depicted by the middle panel of Fig. 2(b). The oscillation pattern of observables is akin to the coupled pendulum model [44], with the two distinct atomic coherences acting as the harmonic oscillators. This is distinct from the limit cycle in purely dissipative cases, e.g., the Lotka–Volterra predator–prey model [45, 46, 47]. Such an understanding is confirmed by adiabatic elimination of coherence terms: in their absence, the system ceases to oscillate.
Figure 2: (a) Mean-field phase diagram in the $\chi$-$\gamma$ plane. (b) Classical trajectories of the system in the $s_{z}$-$n$ plane. The three panels correspond to the three data points indicated in (a). The system is initially in a product state $\rho_{0}=\prod_{i}|{0}\rangle_{i}\langle{0}|$.
As we show later, the observed limit-cycle oscillation is the generic manifestation of a BTC. Such a BTC is not only robust against local decays, but also intrinsically induced by such terms, as indicated by the phase diagram in the $\chi$-$\gamma$ plane [see Fig. 2(a)]. At large decay rate $\gamma$, the system stays in the stationary phase. Here, the strong dissipation suppresses populating the interacting states $\ket{\pm}$, weakening the nonlinearity while relaxing the system to the fixed point [see the left panel of Fig. 2(b)]. As $\gamma$ becomes smaller, a robust limit cycle can be established and maintained for a finite range of $\gamma$, constituting a BTC phase. However, when $\gamma$ further decreases, the nonlinearity dominates over the dissipation and makes the system unstable, which eventually triggers a chaotic motion lacking a well-defined oscillation periodicity [see the right panel Fig. 2(b)]. Conceptually, this makes the observed BTC an intrinsic dissipative time crystal [48], since for a superradiance-based BTC, the ordered oscillation remains in the absence of dissipation.
Full quantum analysis.—Having explored the classical limit of the model, we now turn to a full quantum description. To this end, we first assume an all-to-all ($\alpha=0$) interaction $V_{ij}=\chi/(N-1)$, which is normalized by the system size $N$ to ensure that the interaction energy is extensive. In this regime, the system possesses a weak permutation symmetry $P_{ij}\mathcal{L}P_{ij}^{-1}=\mathcal{L}$ with respect to the permutation $P_{ij}$ between any pair of particles at sites $i$ and $j$ in a mixed state [49, 50]. Within the symmetric subspace, the complexity of the many-body problem reduces from $\mathcal{O}(9^{N})$ to $\mathcal{O}(N^{8})$ [37], allowing us to diagonalize the Liouvillian up to $N\sim 20$. While this special regime has a classical correspondence for separable initial states [51] in the thermodynamic limit ($N\rightarrow\infty$), the system is in general strongly correlated and cannot be captured by the mean-field description.
We first study the Liouvillian spectrum $\{\lambda_{k}\}$ with $\mathcal{L}(\rho_{k})=\lambda_{k}\rho_{k}$ in the BTC phase. As the system size increases from 10 to 20, we identify multiple ordered branches of eigenvalues $\lambda_{k}$ shifting towards the imaginary axis, which is a universal feature for systems supporting limit cycles [52]. The inset of Fig. 3(a) depicts the low-lying spectrum for $N=20$, where we can identify a branch of eigenvalues evenly spaced along the imaginary-axis direction, as indicated by the blue dots close to the dashed grids $0,\pm\omega,\pm 2\omega,\cdots$, with $\omega$ the oscillation frequency of the limit cycle. This branch will collapse onto the imaginary axis in the thermodynamic limit, responsible for all possible oscillation patterns of the BTC. The real part of such a gapless branch follows an algebraic decay $\mathrm{Re}[\lambda_{k}]\propto-N^{-\beta_{k}}$. Importantly, smaller frequency components approach the imaginary axis with a higher rate $\beta_{k}$ [see Fig. 3(a) for the gap closing process of ±$\pm\omega$ and ±$\pm 2\omega$ components]. As a result, the gapless branch will gradually form a parabolic shape, responsible for the quantum state diffusion in the phase space [52].
Figure 3: Characterization of the BTC with all-to-all interactions. (a) Evolution of several low-lying Liouvillian eigenvalues $\lambda_{k}$ with $N$ along the real axis. The data are fitted by an algebraic function $c_{k}+N^{-\beta_{k}}$. The inset shows the spectrum in the complex plane for $N=20$. (b) The two-time correlation function $G(t)$ for different system sizes.
In addition to ordered branches, we find eigenvalues whose imaginary parts are incommensurate with $\omega$. The ones closest to the imaginary axis are indicated by the red diamonds in Fig. 3(a). These Liouvillian eigenstates are gapped, and determine the time scale at which a stable oscillation can be established. The gap $\approx 0.594\gamma$ drawn from an algebraic fitting shows a nice agreement with the decay rate $\approx 0.602\gamma$ toward the limit cycle.
While the spectrum analysis shows a nice quantum-classical correspondence, we now investigate whether the phase can be defined as a BTC. For this, we focus on properties of the unique steady state $\rho_{\mathrm{ss}}$ for a finite $N$ satisfying $\mathcal{L}(\rho_{\mathrm{ss}})=0$, in analogy to the ground state of a closed many-body system. As pointed out in Ref. [8], a continuous time crystal can be unambiguously identified by the behavior of the two-time correlation function. For the considered open system, it is defined as
$G(t)=\mathrm{Tr}[\tilde{S}_{z}(t)\tilde{S}_{z}(0)\rho_{\mathrm{ss}}]/N^{2},$ (3)
where $\tilde{S}_{z}=S_{z}-\langle S_{z}\rangle_{\mathrm{ss}}$ with $S_{z}=\sum_{i}s_{i}^{z}$ an extensive operator having a steady-state value $\langle S_{z}\rangle_{\mathrm{ss}}=\mathrm{Tr}[S_{z}\rho_{\mathrm{ss}}]$. A well-defined BTC should possess a finite and periodic correlation function $G(t)$ as $N\rightarrow\infty$ [11, 53]. Physically, $G(t)\neq 0$ implies that a finite perturbation can trigger a persistent and global oscillation, which is the key feature of a time crystal. Applying the quantum regression theorem [54], we obtain an increasingly persistent $G(t)$ as $N$ grows [see Fig. 3(b)]. The tail of the correlation function can be fitted into the form $\mathrm{Re}[G(t)]=Ae^{-\kappa t}\cos(\omega t+\phi)$, where $\kappa\sim N^{-0.6}$ matches the evolution of the smallest gap, while $A$ converges to a finite value, suggesting a truely persistent oscillation in the thermodynamic limit.
Quantum fluctuations and non-Gaussianity.—From the above calculation, we notice that the equal-time correlation $G(0)$ approaches a nonzero constant, implying that the steady-state fluctuations of $S_{z}$ [55, 56, 57]
$F^{2}=(\langle S_{z}^{2}\rangle_{\mathrm{ss}}-\langle S_{z}\rangle_{\mathrm{ss}}^{2})/N=G(0)N,$ (4)
grows linearly with the system size. Such a divergent behavior of the quantum fluctuations is a universal feature of the BTC, as was also observed in previous studies [55]. The linear scaling between $F^{2}$ and $N$ is verified by a larger scale simulation up to $N=80$ [see Fig. 4(a)], which is obtained by the Monte-Carlo wavefunction based matrix product state (MPS) simulation [58]. For comparison, we also carry out the simulation for a stationary phase [see Fig. 4(b)]. In this regime, $F^{2}$ decreases with the system size and converges to a finite value, in stark contrast to the BTC phase.
Figure 4: Scaling of the steady-state fluctuations $F^{2}$ with $N$. (a) and (b) show results for the BTC phase ($\Delta=-1$) and the SP ($\Delta=-7$), respectively. The insets show the time evolution of the fluctuations ($N=80$ for the BTC and $N=40$ for the SP), obtained from the MPS (red curves) and the CME (blue curves). The errorbars indicate the temporal fluctuations caused by finite trajectory realizations.
We next investigate the non-Gaussianity of the system by applying the cumulant expansion (CME) [37]. Up to the second order, CME assumes a Gaussian-type correlation between particles, i.e, $\langle{o}_{i}{o}_{j}{o}_{k}\rangle=\langle{o}_{i}{o}_{j}\rangle\langle{o}_{k}\rangle+\langle{o}_{i}{o}_{k}\rangle\langle{o}_{j}\rangle+\langle{o}_{j}{o}_{k}\rangle\langle{o}_{i}\rangle-2\langle{o}_{i}\rangle\langle{o}_{j}\rangle\langle{o}_{k}\rangle$ for operators $({o}_{i},{o}_{j},{o}_{k})$ at three different sites. In the stationary phase, the state of the system remains largely Gaussian during the entire evolution, revealed by the good agreement between the CME and the exact/MPS simulations [see the inset of Fig. 4(b)]. However, entering the BTC phase, the discrepancies between the CME and the exact/MPS simulations become increasingly visible with a growing $N$. This implies that the evolution in the BTC phase is generally a complex problem involving intricate many-body correlations, in accordance with a recent work demonstrating the non-stabilizerness of the BTC [59].
Classical to quantum BTC.—Despite the breakdown of exact Gaussian-type correlations, we note that the cumulant expansion can still qualitatively distinguish between the distinct scaling properties of the BTC phase and the stationary phase. We now employ this approach to investigate existence of the BTC under a power-law interaction $V_{ij}\propto|i-j|^{-\alpha}$. To ensure that the system has a constant energy per site, we apply the Kac factor to the interaction strength [60], i.e., $C=\chi/\sum_{r=1}^{N-1}{r^{-\alpha}}$. The drastically reduced complexity $\mathcal{O}(N)$ of the CME (using the translation invariance [37]) allows us to extract the phase diagram shown in Fig. 5(a), which will be elucidated below.
We first investigate the lifetime $\tau$  [^1] The lifetime $\tau$ is defined as the time at which the oscillation contrast reaches a given small value $\epsilon$. While precise values of $\tau$ depend on the specific choice of $\epsilon$, different choices yield the same qualitative results described in the main text, which uses $\epsilon=10^{-5}$. of the oscillating dynamics starting from a separable state $\rho_{0}=\prod_{i}|{0}\rangle_{i}\langle{0}|$. Figure 5(a) shows $\tau$ as a function of $\alpha$. For a small $\alpha$, as $N$ grows from $500$ to $1500$, the lifetime gets prolonged, suggesting a closing Liouvillian gap and a BTC phase, as in the $\alpha=0$ case. However, if the interaction is too short-ranged, $\tau$ converges to a finite value, indicative of a gapped, stationary phase.
The transition from the BTC to the SP occurs abruptly at an $\alpha$ between 1 and 1.5. To study the critical behavior driven by the long-range exponent $\alpha$, we then focus on the long-range off-diagonal correlation in the steady state
$\mathcal{C}_{\mathrm{off}}=\frac{1}{N(N-1)}\sum_{i\neq j}\left[\langle s_{i}^{z}s_{j}^{z}\rangle_{\mathrm{ss}}-\langle s_{i}^{z}\rangle_{\mathrm{ss}}\langle s_{j}^{z}\rangle_{\mathrm{ss}}\right],$ (5)
In fact, a time crystal defined by Eq. (3) possesses not only temporal order but also long-range spatial order, which is indeed true for all-to-all interaction since $\mathcal{C}_{\mathrm{off}}\approx F^{2}/N=G(0)$ remains finite in the large-$N$ limit. As shown in Fig. 5(c), $\mathcal{C}_{\mathrm{off}}$ decreases monotonically with an increasing $\alpha$ and eventually vanishes when entering the SP  [^2] This trend is verified by the MPS for $N=20$ [37]. The critical property is then analyzed through finite-size scaling by assuming a universal form of $\mathcal{C}_{\mathrm{off}}=N^{\zeta/\nu}\mathcal{G}[N^{1/\nu}(\alpha-\alpha_{c})]$. We obtain a nice data collapse [see the inset of Fig. 5(c)] and extract a critical value $\alpha_{c}\approx 1.22$ together with two critical exponents $(\zeta,\nu)\approx(-0.56,1.40)$. We note that $\alpha_{c}$ is generally larger than 1 but not rigid under variation of system parameters, typical for long-range interacting systems [63].
A critical $\alpha_{c}$ larger than the system dimension allows for a further classification within the BTC phase. For an uncorrelated initial state, $\alpha\leq 1$ yields a locally noninteracting system as $N\rightarrow\infty$ [51], such that evolution of local observables can be predicted by the mean-field theory. In this regime, the oscillation can be interpreted as a classical limit cycle, representing a classical BTC. However, for $1<\alpha<\alpha_{c}$, the interaction strength $V_{ij}$ between two arbitrary particles remains finite in the thermodynamic limit, facilitating establishment of spatial correlations at finite time. The correlation subsequently prevents a mean-field description of the oscillation, which, hence can be dubbed as a quantum BTC. To demonstrate this, we show in Fig. 5(d) the mean correlation $\mathcal{C}_{\mathrm{off}}^{r}=\sum_{d=1}^{r}[\langle s_{i}^{z}s_{i+d}^{z}\rangle-\langle s_{i}^{z}\rangle\langle s_{i+d}^{z}\rangle]/r$ within a given range $r$ during the oscillation. As $N$ increases, $\mathcal{C}_{\mathrm{off}}^{r}$ becomes vanishingly small for $0\leq\alpha\leq 1$, but is still nonzero for $\alpha>1$. We remark that the notion of classical and quantum BTCs is based on a dynamical classification, i.e., if the initial state $\rho_{0}$ carries correlations (cannot be factorized), a BTC with $\alpha\leq 1$ does not have a classical correspondence, either.
Figure 5: (a) A phase transition driven by the factor $\alpha$. (b)-(d) show the lifetime $\tau$, the steady-state off-diagonal correlation $\mathcal{C}_{\mathrm{off}}$, and the mean correlation $\mathcal{C}_{\mathrm{off}}^{r}$ ($r=30$ and averaged over one period at $t=30$) as a function of $\alpha$ for the indicated system sizes ($N=500,1000,1500$). The results are obtained from the CME. (e) A single-trajectory MPS simulation of the dynamics for $\alpha=1.1$ (green) and $1.6$ (grey), whose Fourier spectra are shown in the inset.
The existence of a quantum BTC is verified by a single-trajectory MPS simulation, which is shown to be capable of unveiling a BTC phase for large system size [64]. Figure 5(e) displays the MPS simulations with power-law interactions [65] for $N=100$. In the predicted stationary phase ($\alpha=1.6$), we observe a disordered oscillation of $s_{z}$, mainly caused by the random quantum jumps. In contrast, the oscillation in the quantum BTC regime ($\alpha=1.1$) has an obviously larger contrast, as well as a more ordered structure supported by a much more pronounced peak in the Fourier spectrum (see the inset).
Conclusions and outlook.—To summarize, we investigate the dynamical phase of an open many-body system composed of long-range interacting spin-1 particles. We demonstrate existence of the BTC in a wide parameter range, supported by an extensive analysis involving mean-field dynamics, Liouvillian spectrum, as well as temporal and spatial correlation functions. Such a highly nontrivial phase survives in the presence of local decay, making it fundamentally more robust than previously discovered BTCs. Moreover, we show that by decreasing the range of the interaction, a BTC that has a classical correspondence undergoes a transition into a quantum version, which carries nonvanishing correlations and cannot be interpreted as a limit cycle.
The proposed model can be potentially realized in artificial quantum systems, such as trapped ions or neutral atoms. In the former case, tunable power-law interactions with $30\leq\alpha<$ 3 can be achieved [66, 67, 68], aligning with the range considered in this work. For the latter, all-to-all interactions can be induced by cavity QED [69, 70] or Rydberg dressing in atomic ensembles [71]. Our model can also be generalized to higher dimensions, where the long-time dynamics may be accessed by improved Krylov methods [72]. In higher dimensions, dipolar ($\alpha=3$) or van der Waals ($\alpha=6$) interactions offered by Rydberg-atom arrays [73, 74, 75] may be sufficiently long-ranged to stabilize the BTC. In addition, our studies may reveal a new universality class for nonequilibrium phases [76], and open the door to exploring versatile dynamical phases, such as multistability [77, 78, 79] and measurement-induced phase transition [80].
Acknowledgements.
Acknowledgments.—We thank T. Pohl, X.L. Li, Y. Zhang, G.M. Zhang for providing helpful suggestions, and X.Y. Wen, Z. Fu, C.P. Qin for sharing computation resources. F.Y. acknowledge valuable discussions with T. Vovk, L. Sieberer, and H. Pichler. This work is supported by the Innovation Program for Quantum Science and Technology (2021ZD0302100) and by the NSFC (Grants No. 12361131576, No. 92265205). B.B. acknowledges the support from funding by the French National Research Agency (ANR) under project ANR-24-CPJ1-0150-01 and a research grant (42085) from VILLUM FONDEN. K.M. acknowledges the support from the Carlsberg Semper Ardens project QCooL. Work in Innsbruck is supported by the ERC Starting grant QARA (Grant No. 101041435). Note added: During preparation of this manuscript, we became aware of a related work on three-level dissipative time crystals induced by a dipolar interaction in 2D [81].